Show that the line x + y − 2 = 0 is a tangent to the parabola x2+ 8y = 0 and find the position vector of the point of contact.


1
Expert's answer
2021-08-09T14:51:08-0400
"x^2+8y=0"

"y=-\dfrac{1}{8}x^2"

"y'=-\dfrac{1}{4}x"

"y-y_1=y'(x-x_1)"

"y=-\dfrac{1}{4}x_1(x-x_1)-\dfrac{1}{8}x_1^2"

"y=-\dfrac{1}{8}x_1x+\dfrac{1}{4}x_1^2"

"x + y \u2212 2 = 0=>y=-x+2"

"\begin{matrix}
 -\dfrac{1}{4}x_1=-1 \\
\\
 \dfrac{1}{8}x_1^2=2
\end{matrix}"

The system has the solution "x_1=4."


"y_1=-\dfrac{1}{8}(4)^2=-2"

The line "x+y-2=0" is a tangent to the parabola "x^2+8y=0" at the point "(4, -2)."

The position vector of the point of contact is "\langle4, -2\rangle."